# Theorem-Carrying-Transaction: Runtime Certification to Ensure Safety for Smart Contract Transactions

### v1

@@ 5.3 Concolic executor
Figure 7: Concrete execution vs. symbolic execution.
Example. The concolic executor requires substantial understanding about the Solidity specification [18] and the EVM specification [17]. Due to space constraints, we skip many details and use the EVM code segment in Figure 8 as an example. The figure shows the initial stack and the stack before 925 JUMPI. Essentially, the code pushes two elements onto the stack top: the first is 0x03a3 (i.e., decimal value 931); the second is ISZERO (LT (SLOAD (MapElement (MultiVulnToken.balances, Partial32B((12, 31), _from))), EVMADD(_fee, _value))).
Figure 8: Execution of an EVM code segment.
A thorough understanding of the EVM code is not required for this section, but some clarifications may help.
  • The EVM storage is a key-value pair. Both the key and the value are 32-byte. The key is obtained by computing a SHA3 over the slot number of the map variable and the index of the element according to reference [22]. In our example, the slot number is 2 (i.e., 878 PUSH1 02), which corresponds to the map variable balances based on the compilation symbol info of MultiVulnToken. The index is Partial32B((12,31), _from). Hence, the MapElement subtree is equivalent to balances[_from] at the Solidity level.
Boogie code generation. When the concolic executor gets to 925 JUMPI, which is a branch, it calls the code generation functionality (a.k.a. CodeGen). The fact that the next instruction is 931 (i.e., the first stack element 0x03a3) indicates that the branch is taken, so the branch condition (i.e., the second stack element) should be assumed true. CodeGen performs a post-order traversal of the SVT to generate the following Boogie code. Note that Boogie does not use the object-oriented (OO) notation “obj.var” to represent a variable in an object. Instead, it is modeled as a dictionary item ClassName.var[objRef], as described in reference [11]. Thus, the first line would mean entry_contract.balances[_from] if an OO notation was used. With 7, we can see that the EVM code segment in Figure 8 is equivalent to the Solidity statement require (balances[_from] >= _fee + _value) in Figure 2.
@@
tmp4 := !tmp3;
assume tmp4;
Code 7: The Boogie code generated from Figure 8.
Inter-contract calls. A significant advantage of TCT’s concolic approach is the easiness to handle a transaction involving multiple contracts. The reentrancy transaction in Section 3 is a simple example. In Section 6, we will show that a Uniswap transaction usually involves five contracts. These contracts call each other through address-references, so it is challenging for a static code analyzer to resolve the addresses of all potential callees. TCT does not face this challenge, as every call is resolved in the concrete trace.
@@
We implement three components to support inter-contract calls. First, we give every contract its own stack and memory like in a classical EVM. Second, we follow the Ethereum yellow paper [17] to implement call instructions (e.g., CALL, STATICCALL) and return instructions (e.g., RETURN, STOP). This requires correctly modeling EVM’s marshalling and unmarshalling mechanisms for call arguments. Third, the call stack also needs to hold symbolic values, because a callee contract address is often a result of the caller contract’s computation. Like the data stack in Figure 8, we implement the call stack so that its elements are also SVT references.
The entire multi-contract transaction is generated as a single Boogie function, in which every contract variable is fully qualified by its contract. For example, Multi VulnToken.balances[entry_contract][_from] is the Boogie representation of balances[_from] of the contract at the bottom of the call stack, i.e., entry_contract.
5.4 Concrete checking by Geth
We have described how theorem proving is prototyped in TCT. Whether a theorem is applicable to a transaction requires concretely checking hypothesis $\varphi(v,s)$ and computing a path hash $ph$. We prototype these in Geth 1.13.0.
@@ 5.4 Concrete checking by Geth
Hypothesis-checking code is hooked into the beginning of function applyTransaction in the Geth core. We follow the Solidity documentation [22] to get the layout of variables in a contract. We use the Geth function StateDB::GetState to read contract variables.
@@
To compute the path hash, we add a byte buffer to the Interpreter struct in the EVM module. We then modify the file instructions.go so that a branch instruction (i.e., JUMPI) records the target program counter, and a call instruction (i.e., CALL, STATICCALL, DELEGATECALL, CALLCODE) records the callee address. The path hash, which is the Keccak-256 hash of the buffer, is computed at the end of function applyTransaction.
@@
As TCT’s name suggests, every transaction carries a theorem. This is implemented by adding a theorem_hash field, described in the last paragraph in Section 4, in the Transaction struct in the Geth core.
6 Scaling up from ERC20 to Uniswap
@@
ERC20 token is a major application on Ethereum. Section 3 shows one example. To show TCT’s generality, we have applied it to the transfer transactions of the top-5 ERC20 contracts according to Etherscan [23], namely USDT, BNB, USDC, stETH and TONCOIN. These transfers are not more complex than what we show in Section 3. The observation supports our premise – although smart contract’s static code can be complex, each transaction is usually short and simple for symbolic verification, thus TCT can scale to sophisticated cross-contract scenarios.
To test our approach on scenarios that are significantly more complex, we apply TCT to a codebase of the Uniswap decentralized exchange. Uniswap is the most extensively used smart contract application on Ethereum based on the amount of gas burnt [24] (as of May 2024). We now present a very concise explanation about the design of Uniswap’s basic mechanism, based on which we explain how TCT is applied.
6.1 Concise explanation of Uniswap’s design
@@
Uniswap’s core design is an automatic market maker (AMM) algorithm to allow users holding different ERC20 tokens to form a market, which dynamically and automatically sets prices for these tokens. Assuming most users want to maximize their gains from the trades, the prices will reach an equilibrium that represents the fair market prices.
Figure 9: The design of Uniswap’s AMM algorithm.
Uniswap has evolved to version V3 [25], with version V4 in development [26]. We study version V2 [27] because it already contains the essentials in the AMM mechanism, whereas V3’s focus is to add a notion of “concentrated liquidity” to augment the AMM, and V4’s focus is to make Uniswap customizable by developers.
@@
The other user type is trader. Traders do not swap tokens with each other, but with the corresponding pair. For example, if traders want to swap TokenA for TokenB or vice versa, they invoke the swap operation with the TokenA-TokenB pair in Figure 9. The AMM algorithm requires the reserve amounts to maintain inverse proportionality, which is also known as the constant-product relation. Suppose the current point is $(x_{1},y_{1})$. A swap operation can move the point to $(x^{\prime}_{1},y^{\prime}_{1})$ along the hyperbola $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$. From the trader’s perspective, they give $\Delta x_{1}=x^{\prime}_{1}-x_{1}$ amount of TokenA to the pair and gets $\Delta y_{1}=y_{1}-y^{\prime}_{1}$ amount of TokenB from the pair. The relative price is $\Delta x_{1}/\Delta y_{1}$. If traders try to profit from trading, the pair’s reserves (i.e., the black dot) will settle in an equilibrium representing the market’s consensus on the relative price between the two tokens. When the market is in action, the operations of LPs and traders interleave, so the dot can move to any point inside the quadrant, e.g., along the path $(x_{1},y_{1})\rightarrow(x_{2},y_{2})\rightarrow(x^{\prime}_{2},y^{\prime}_{2})$.
Safety and fairness. The proportionality and inverse proportionality are the foundation of Uniswap’s design. The founders of Uniswap contracted a company to formally analyze the design and published the analysis report [28, 29]. LP’s proportionality ensures that every LP can get the proportional share of the reserves back at any time. Trader’s inverse proportionality ensures that, if a trader swaps $\Delta x$ TokenA for $\Delta y$ TokenB, anybody can immediately swap $\Delta y$ TokenB to get $\Delta x$ TokenA. We believe that these are the best to say about the algorithm being “safe” and “fair”. For a currency exchange market, we cannot demand “safety” and “fairness” to mean that nobody has any financial loss. However, gains and losses should be the result of investment decisions, rather than unexpected code behaviors.
6.2 Applying TCT on Uniswap
@@
We target the UniswapV2-Solc0.8 implementation [30], which consists of four contracts: the pair contract implements a pair as described above; the UniswapERC20 contract implements the liquidity token for a pair; the factory contract implements the whole market, which creates and stores all pairs; the router contract provides entry functions for the user. Because a liquidity token is specific to a pair, the code derives the pair contract from the UniswapERC20 contract. Thus, an instance of the pair contract behaves both as a reserve of two tradable tokens, and as its own liquidity token.
@@
8 shows the properties defined as postconditions for functions addLiquidity and swap in the router contract. The postcondition of removeLiquidity is identical to that of addLiquidity, so it is not shown. The declaration and assignment lines are used to make the postconditions concise. The operator old is a notation of Boogie. TCT uses it to represent the value of a variable at the beginning of the transaction execution. Note that pair.totalSupply is the amount of liquidity token that this pair contract has minted.
Because these postconditions are defined over three contracts: tokenA, tokenB and their pair (i.e., pair), it is problematic for a static code verifier to examine the safety of the Uniswap’s own contracts (i.e., pair, factory and router) alone, without the concrete code of TokenA and TokenB. For example, a hypothetical TokenA implementation can simply change tokenA.balanceOf[pair] to violate the postconditions, which is the external code totally out of the Uniswap programmers’ control. As we discuss earlier, this is a dilemma for static verification because it needs to make assumptions about external and future implementations.
Our prototype uses a simple token code to construct both TokenA and TokenB. The code is derived from the ERC20 base contract, which specifies the invariants in 6. TCT guarantees that every transaction satisfies the properties of both Uniswap and the tokens.
///@custom:declaration var factory,pair: address;
@@
///@custom:assignment factory:=this.factory; pair:=factory.getPair[tokenA][tokenB];
@@
function addLiquidity(address tokenA, address tokenB, uint256 amountADesired,
uint256 amountBDesired, uint256 amountAMin,
uint256 amountBMin, address to
///@custom:declaration var factory,pair:address;
///@custom:assignment factory:=this.factory; pair:=factory.getPair[tokenA][tokenB];
@@
function swap(
uint256 amountIn, uint256 amountOutMin,
address tokenA, address tokenB, address to
Code 8: The properties for liquidity and swap operations.
6.3 Higher complexity
Uniswap uses more Solidity language features than those in the studied ERC20 contracts. They include loops, libraries, two-dimensional mappings and dynamic arrays. Every liquidity or swap transaction makes inter-contract calls across many contracts. Our concolic executor correctly converts these complex EVM traces into Boogie programs. Section 7 will explain that the hypotheses of the theorems and the underlying mathematics for the proofs are also much more complex.
7 Evaluation
In this section, we provide quantitative measurements and detailed analyses to show the code complexity that the TCT prototype has handled, the runtime overhead it has achieved, and new code insights it has revealed.
7.1 Code complexity
Table I shows the lines-of-code (LoC) numbers of six EVM traces. The first five rows are the tranfer transactions of the top-5 ERC tokens. Next, TrfrProxy corresponds to the transaction calling MultiVulnToken::transferProxy. Rntrcy and NoRntrcy call MultiVulnToken::clear, the former exploiting the reentrancy pattern and the latter does not. The other traces correspond to Uniswap transactions addLiquidity, removeLiquidity and swap, which are much more complex. For example, Swap contains 9495 EVM instructions involving 5 different contracts. It makes 10 cross-contract calls. The concolic executor generates 299 lines of Boogie code, excluding variable declarations and axioms. The appendix shows the Boogie code for TrfrProxy as an example. All other traces are available in a GitHub repo which we will disclose in the future.
TABLE I: Lines of code of the traces.
@@ 7.1 Code complexity
USDTTrfr 904 1 0 72
BNBTrfr 686 1 0 64
USDCTrfr 736 1 0 46
stEthTrfr 1119 1 0 60
TONCOINTrfr 683 1 0 32
TrfrProxy 704 1 0 47
Rntrcy 1385 2 3 52
NoRntrcy 628 2 1 21
AddLqdty 7592 5 10 223
RmvLqdty 6985 5 10 199
Swap 9495 5 10 299
7.2 Complexity of the hypotheses and math
Table II shows the hypotheses for the Uniswap traces. They are much more complex than that of TrfrProxy explained in the previous sections. However, with the understanding about Uniswap code, we believe that none of these individual conditions is counterintuitive to the Uniswap programmers. The TCT approach simply explicates their intentions. Under the hypotheses, the traces go through non-trivial computations, equivalent to the math equations shown in the table. Note that tx.origin is a well-defined EVM value, which represents the address of the transaction issuer, i.e., the msg.sender of the transaction’s entry function.
TABLE II: Hypotheses of AddLqdty, RmvLqdty, Swap.
AddLqdty tokenA != tokenB &&
pair.totalSupply > 0 && tx.origin != pair &&
pair.token1 == tokenA && pair.token0 == tokenB &&
tokenA.balanceOf[pair] > 0 && tokenB.balanceOf[pair]
> 0 && tokenA.totalSupply < $\texttt{2}^{\texttt{255}}$& tokenB.totalSupply
< $\texttt{2}^{\texttt{255}}$& pair.reserve0 == tokenB.balanceOf[pair] &&
pair.reserve1 == tokenA.balanceOf[pair]

### v2

@@ 5.2. Concolic executor
Figure 6. Concrete execution vs. symbolic execution.
Figure 7 shows a code trace segment, as well as the initial stack and the stack before 925 JUMPI. The code pushes two elements onto the stack top: one is 0x03a3 (i.e., jump destination 931); the other is ISZERO (LT (SLOAD (MapElement (MultiVulnToken.balances, Partial32B((12, 31), _from))), EVMADD(_fee, _value))). Note that an Ethereum address is 20 bytes, which is denoted by the node Partial32B((12, 31), ...), i.e., the lower 20 bytes of a 32-byte value. More details are given in Appendix A.3, which are not necessary for the rest of this section.
@@
assume !tmp3;
Code 6: The Boogie code generated from Figure 7.
Boogie code generation. When the concolic executor gets to 925 JUMPI, which is a branch, it calls the code generation functionality (a.k.a. CodeGen). The fact that the next instruction is not 926, but 931 (i.e., the first stack element 0x03a3) indicates that the branch is taken, so the branch condition (i.e., the second stack element) should be assumed true. CodeGen performs a post-order traversal of the SVT to generate the Boogie code in LABEL:code:boogieexpl. Note that Boogie does not use the object-oriented (OO) notation “obj.var” to represent a variable in an object. Instead, it is modeled as a dictionary item ClassName.var[objRef], as described in reference (leino_boogie_2, 13). Thus, the first line would mean entry_contract.balances[_from] if an OO notation was used. With LABEL:code:boogieexpl, we can see that the EVM code segment in Figure 7 is exactly the Solidity statement require (balances[_from] >= _fee + _value) in Figure 3.
Figure 7. Execution of an EVM code trace segment.
It is important to note that our task is not “de-compilation”. It is not to reverse-engineer high-level code from low-level instructions using heuristics. As explained earlier, we have the Solidity source code and complete symbol info. The SVT-based Boogie code generation is a deterministic algorithm without heuristic guessing.
For a function postcondition, a Boogie assert is added at the return point of the function in the trace. For a contract invariant, an asssume is added at the entry and an assert is added at the exit of the contract.
@@
Inter-contract calls. A crucial ability of TCT is to handle a transaction involving multiple contracts. The reentrancy transaction in Section 3 is a simple example. In Section 6, we will show that a Uniswap transaction usually involves five contracts. We implement three components to support inter-contract calls. First, we give every contract its own stack and memory like in a classical EVM. Second, we follow the Ethereum yellow paper (wood2014ethereum, 39) to implement call instructions (e.g., CALL, STATICCALL) and return instructions (e.g., RETURN, STOP). This requires correctly modeling EVM’s marshalling and unmarshalling mechanisms for call arguments. Third, the call stack also needs to hold symbolic values, because a callee contract address is often a result of the caller contract’s computation. Like the data stack in Figure 7, we implement the call stack which stores every element as a SVT reference as well.
5.3. Runtime mechanisms added to Geth
@@
In the TCT protocol, whether a theorem is applicable to a transaction requires concretely checking guard $\varphi(s,f.p)$ and computing a path hash $ph$. We implemented the mechanism in Geth 1.13.0. The checking code is hooked into the beginning of function applyTrans- action in the Geth core. We follow the Solidity documentation (soliditylayout, 29) to get the layout of variables in a contract. We use the Geth function StateDB::GetState to read contract variables.
@@
To compute the path hash, we add a byte buffer to the Interpreter struct in the EVM module. We then modify the file instructions.go so that a branch instruction (i.e., JUMPI) records the target program counter, and a call instruction (e.g., CALL, STATICCALL) records the callee address. The path hash, which is the Keccak-256 hash of the buffer, is computed at the end of function applyTransaction.
@@
For a TX to carry theorem $\tau$, we add a theorem_hash field in the Transaction struct in the Geth core. The hash is described in the last paragraph in Section 4.
6. Scaling up from ERC20 to Uniswap
@@
ERC20 token is a major application on Ethereum, and thus represents the classical territory for existing verification techniques. To show TCT’s generality, we have applied it to the transfer transactions of the top-5 ERC20 contracts according to Etherscan (top-tokens, 6), namely USDT, BNB, USDC, stETH and TONCOIN. These transfers are not more complex than what we show in Section 3, and are representative of most ERC20 contracts. The observation supports our premise – although the static code of the GP is gigantic, each transaction at runtime is usually short and simple, which gives TCT a unique advantage to scale to sophisticated scenarios.
To leapfrog from the classical territory to a level of significant complexity, we apply TCT to a codebase of the Uniswap decentralized exchange. Uniswap is the most extensively used DeFi application on Ethereum based on the amount of gas burnt (ultrasoundmoney, 34). Uniswap is at the high end of complexity, using more Solidity language features than those in the studied ERC20 contracts, including loops, libraries, two-dimensional mappings and dynamic arrays. Every liquidity or swap transaction makes inter-contract calls across many contracts. Our concolic executor correctly converts these complex EVM traces into Boogie programs. Section 7 will explain that the theorems and the underlying mathematics for the proofs are more complex.
6.1. Concise explanation of Uniswap’s design
We now give a very concise explanation about Uniswap’s basic mechanism, based on which we define the interface specifications.
@@ 6.1. Concise explanation of Uniswap’s design
Uniswap’s core design is an automatic market maker (AMM) algorithm to allow users holding different ERC20 tokens to form a market, which dynamically and automatically sets prices for these tokens. Assuming most users want to maximize their gains from the trades, the prices will reach an equilibrium that represents the fair market prices. Uniswap has evolved to version V3 (uniswap-v3-protocol, 37), with version V4 in development (uniswap-v4-protocol, 35). We study version V2 (uniswap-v2-protocol, 3) because it already contains the essentials in the AMM mechanism, whereas V3’s focus is to add a notion of “concentrated liquidity” to augment the AMM, and V4’s focus is to make Uniswap customizable.
@@
The other user type is trader. Traders do not swap tokens with each other, but with the corresponding pair. For example, if traders want to swap TokenA for TokenB or vice versa, they invoke the swap operation with the TokenA-TokenB pair in Figure 8. The AMM algorithm requires the reserve amounts to maintain inverse proportionality, which is also known as the constant-product relation. Suppose the current point is $(x_{1},y_{1})$. A swap operation can move the point to $(x^{\prime}_{1},y^{\prime}_{1})$ along the hyperbola $x_{1}\times y_{1}=x^{\prime}_{1}\times y^{\prime}_{1}$. From the trader’s perspective, they give $\Delta x_{1}=x^{\prime}_{1}-x_{1}$ amount of TokenA to the pair and gets $\Delta y_{1}=y_{1}-y^{\prime}_{1}$ amount of TokenB from the pair. The relative price is $\Delta x_{1}/\Delta y_{1}$. If all traders try to profit from trading, the pair’s reserve (i.e., the black dot) will settle in an equilibrium representing the market’s consensus on the relative price between the two tokens. When the market is in action, the operations of LPs and traders interleave, so the dot can move to any point inside the quadrant, e.g., along the path $(x_{1},y_{1})\rightarrow(x_{2},y_{2})\rightarrow(x^{\prime}_{2},y^{\prime}_{2})$.
The proportionality and inverse proportionality are the foundation of Uniswap’s design. The founders of Uniswap contracted a company to formally analyze the mathematical design and published the analysis report (uniswap_blog, 36, 41). The analysis did not contain code verification, which we perform in this paper.
6.2. Applying TCT on Uniswap
@@
We target the UniswapV2-Solc0.8 implementation (uniswapsol, 21), which consists of four contracts: the pair contract implements a token-pair as described above; the UniswapERC20 contract implements the liquidity token for a pair; the factory contract implements the whole market, which creates and stores all pairs; the router contract provides entry functions for the user. Because a liquidity token is specific to a pair, the code derives the pair contract from the UniswapERC20 contract. Thus, an instance of the pair contract behaves both as a reserve of two tradable tokens, and as its own liquidity token.
@@
LABEL:code:prop shows the properties defined as postconditions for functions addLiquidity and swap in the router contract. The postcondition of removeLiquidity is identical to that of addLiquidity, so it is not shown. The declaration and assignment lines are used to make the postconditions concise. The operator old is a notation of Boogie. TCT uses it to represent the value of a variable at the beginning of the transaction execution. Note that pair.totalSupply is the amount of liquidity token that this pair contract has minted.
///@custom:declaration var pair: address;
@@
///@custom:assignment pair:=this.factory.getPair[tokenA][tokenB];
@@
function addLiquidity(address tokenA, address tokenB, uint256 amountADesired, uint256 amountBDesired, uint256 amountAMin, uint256 amountBMin, address to
///@custom:declaration var pair:address;
///@custom:assignment pair:=this.factory.getPair[tokenA][tokenB];
@@
function swap(uint256 amountIn, uint256 amountOutMin, address tokenA, address tokenB, address to) { ... }
Code 7: The properties for liquidity and swap operations.
The dilemma for static verification. Although these postconditions are defined for Uniswap functions, they are across three external contracts: tokenA, tokenB and pair (i.e., the pair-contract to swap TokenA and TokenB). The same dilemma discussed in Section 3 also exists if we ask “whether the code of Uniswap (i.e., pair, factory and router alone) satisfies the proportionality and inverse-proportionality”. The answer depends on the concrete code of the specific token contracts out of the Uniswap programmers’ control. Even reading a state variable inside an external contract relies on the implementation of a getter function in the contract. Thus, the question is only meaningful w.r.t. each and every transaction that operates concrete tokenA and tokenB contracts at runtime.
7. Evaluation
We now analyze and measure TCT’s practicality and performance.
7.1. Branches and loops in real-world contracts
As explained in Section 2.2, although there are millions of contracts in the $GP$, the complexity of each contract is quite low, meaning that a small number of traces are sufficient to cover all the behavior of a contract.
We performed a quantitative study to support this observation, analyzing code complexity in terms of the loops and branches found in practice. We download 1760, randomly selected, real transactions on Ethereum between years 2022 and 2024 with the criterion that the entry contract of each transaction makes 3 $\sim$ 10 inter-contract calls, which represents the typical complexity range: we consider a transaction too simple if the entry contract makes fewer than 3 such calls, whereas an entry contract with more than 10 such calls often corresponds to a batch transaction, which bundles several operations that could otherwise be individual transactions. After grouping and de-duplicating similar transactions, the set contains 275 cases. We will make this dataset public.
For each case, we take the entry function. If the function code is less than 5 lines, we include its most significant callee so that the function code is non-trivial. The average lines-of-code (LoC) of these functions, which includes comment lines inside function bodies, is 41.10, shown in Table 1.
Loops. There are 54 functions containing loops. Among them, 52 functions use loops to iterate through an array of contracts (usually token contracts) and apply the same operations to each one. These are batching scenarios: every iteration could be its own transaction, but the loop batches all iterations into one transaction for two main reasons: transactional guarantees and gas saving. In the remaining two cases, loops are simple: one calculates the sum of an integer array; the other initializes every array element to a constant. Note that although the dataset contains entry functions, we also search for loops in the entire contracts, and conclude that loops are significantly more scarce in non-entry functions.
Table 1. Study about smart contract loops and branches.
# of
functions
Avg.
LoC
# of “non-
batching” loops
Avg.
branches
Avg. “harder-to-
merge” branches
275 41.10 2 (out of 54) 1.301 0.298
Branches. Branches include if-statements and question-mark-expressions. However, “if (...) revert” is not a branch, as it is equivalent to a require-statement. The average number of executed branches is 1.301 per function, which already is very low. We further classify branches as follows: (1) for a branch if (...) {B1} else {B2}, if both B1 and B2 make one or more function calls (either inter-contract or cross-contract), we say that it is “harder-to-merge”. Otherwise the branch is “easy-to-merge”; (2) branches without the else-clause are also “easy-to-merge”. The average number of harder-to-merge branches per function is 0.298.
Consider two traces that only diverge at $n$ easy-to-merge branches. It is easy to merge the two traces into one Boogie code (using Boogie’s if-statements), so the theorem of the code will cover $2^{n}$ paths. In other words, the number of needed theorems (traces) do not grow exponentially with the number of easy-to-merge branches. Although our current prototype has not implemented the merging mechanism, the purpose of this study is to show that the complexity of branches is low in real-world contracts.
Table 2. Stats of the transaction traces in our case studies.
@@ 7.1. Branches and loops in real-world contracts
USDTTrfr 904 1 0 0 2 0 72
BNBTrfr 686 1 0 0 0 0 64
USDCTrfr 736 1 0 0 0 0 46
stEthTrfr 1119 1 0 0 0 0 60
TONCOINTrfr 683 1 0 0 0 0 32
TrfrProxy 704 1 0 0 0 0 47
Rntrcy 1385 2 3 0 0 0 52
NoRntrcy 628 2 1 0 0 0 21
AddLqdty 7592 5 10 0 4 1 223
RmvLqdty 6985 5 10 0 2 1 199
Swap 9495 5 10 1 4 1 299
7.2. Transaction traces verified by TCT
Table 2 shows 11 transaction traces. The first five rows are the tranfer transactions of the top-5 ERC tokens. Next, TrfrProxy denotes the transaction calling MultiVulnToken::transferProxy (see Section 3). Rntrcy and NoRntrcy call MultiVulnToken::clear, both containing cross-contract calls. The former exploits the reentrancy pattern and the latter does not (also see Section 3) The last three traces correspond to Uniswap transactions addLiquidity, removeLiquidity and swap, which are much more complex. For example, Swap contains 9495 EVM instructions involving 5 different contracts. It makes 10 cross-contract calls. The Boogie code consists of 299 lines, excluding variable declarations and axioms. The “harder-to-merge” branch is a case-split on the condition address (TokenA) $\leq$ address(TokenB). From the perspective of the TokenA-TokenB-pair contract, swapping TokenA for TokenB and swapping TokenB for TokenA are dual cases. The loop is used for batching.
7.3. Complexity of the guards and math
Table 3 shows the guards for the Uniswap traces. They are much more complex than that of TrfrProxy explained earlier. However, none of these individual conditions should be counterintuitive to the Uniswap programmers. The TCT approach simply explicates their intentions. Under the guards, the traces go through non-trivial computations, equivalent to the math equations shown in the table. Note that tx.origin is a well-defined EVM variable, which represents the address of the transaction issuer, i.e., the msg.sender of the transaction’s entry function.
Table 3. Guards for AddLqdty, RmvLqdty and Swap, as well as the underlying math computations.
AddLqdty tokenA != tokenB && pair.totalSupply > 0 && tx.
origin != pair && pair.token1 == tokenA && pair.
token0 == tokenB && tokenA.balanceOf[pair] > 0 &&
tokenB.balanceOf[pair] > 0 && tokenA.totalSupply <
$\texttt{2}^{\texttt{255}}$ && tokenB.totalSupply < $\texttt{2}^{\texttt{255}}$ && pair.reserve0
== tokenB.balanceOf[pair] && pair.reserve1
== tokenA.balanceOf[pair]